OpenStudy (anonymous):

\[\huge \dfrac{\sqrt{2}}{\sqrt[3]{2}} = ?\]

ganeshie8 (ganeshie8):

\[\huge \dfrac{\sqrt{2}}{\sqrt[3]{2}}\]

OpenStudy (anonymous):

\[\huge a^na^m=a ^{n+m}\]

OpenStudy (anonymous):

If dividing subtract the exponents.

OpenStudy (anonymous):

\[\huge \sqrt[n]{x^m}\implies x ^{\frac{ m }{ n }}\]

OpenStudy (midhun.madhu1987):